CAVITY RINGDOWN ABSORPTION SPECTRUM OF THE T_{1}(n, π^{∗}) ← S_{0} TRANSITION OF 4-CYCLOPENTEN-1,3-DIONE

M. G. SPRINGER, S. P. JAGUSCH, N. C. HLAVACEK, and \underline{S. DRUCKER}, Department of Chemistry, University of Wisconsin-Eau Claire, Eau Claire, WI 54702.

The cavity ringdown absorption spectrum of 4-cyclopenten-1,3-dione (CPD) was recorded near 487 nm in a roomtemperature gas cell. The very weak band system (ϵ < 0.05 dm^{3} mol^{−1} cm^{−1}) in this region is due to the T_{1}(n, π*) ← S_{0} electronic transition. The origin band was observed at 20,541±1 cm^{−1}. We have assigned 15 vibronic transitions in a region extending to about +700 cm^{−1} relative to the origin band. From these tentative assignments we determined fundamental frequencies for several vibrational modes in the T_{1} excited state. These include the out-of-plane ring modes ν′_{19} (b_{1}) and ν′_{14} (a_{2}). The table below compares their frequencies to corresponding values^{a} in the S_{0} electronic ground state and the S_{1}(n, π*) excited state.

Ring frequencies (cm^{−1}) of CPD in its lowest electronic states

Mode Description S_{0} S_{1} T_{1}
ν_{19}  pucker  99 160 106
ν_{14}  twist  239 307 299

The increases in these ring frequencies upon electronic excitation signify that the nominal n → π^{∗} chromophore is delocalized to include the conjugated ring atoms. It is noteworthy, though, that the ν_{19} fundamental increases by only 7 cm^{−1} upon T_{1} ← S_{0} excitation, compared to 61 cm^{−1} for the S_{1} excitation. Other cyclic conjugated enones show differences (S_{1} vs. T_{1} ring frequencies) of similar magnitude. These findings are attributable to the configuration mixing that each of the excited states undergoes. The two n, π^{∗} excited states may undergo mixing within distinctly different manifolds of spin-orbitals.

aR. A. Back and R. D. Gordon, J. Molec. Spectrosc. 204, 85 (2000).